I need help on this promblem on solving equations.
2x+75=102-x

i tried 2 solve it but i think its wrong.
2x+75=102-x
2x+x=75-102
3x=-27
is it right???

Answer by rapaljer(4671)

You are correct down to the next to last step.
You have this: 2x+x=75-102

It should be:
2x+x=102-75
3x=27
x=9

R^2 at SCC